40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 7 19:42:01 2018, gautam, Update, General, IFO checkup and DRMI locking prep 
    Reply  Wed Nov 7 23:07:11 2018, gautam, Update, LSC, DRMI locking recovered Screenshot_from_2018-11-07_23-05-58.pngSRCL2MICH_crosscpl.pdfPRCangularCoh_rot.pdf
       Reply  Sat Nov 10 17:32:33 2018, gautam, Update, LSC, Nulling MICH->PRCL and MICH->SRCL sensMat_2018-11-10.pdf
          Reply  Tue Nov 13 18:09:24 2018, gautam, Update, LSC, Investigation of SRCL-->MICH coupling TF_sweptSineMeas.pdfdigitalDemod.pdfdigitalDemod_hist.pdfDRMI_LSCspectrogram.pdf
             Reply  Fri Nov 16 00:47:43 2018, gautam, Update, LSC, More DRMI characterization PRCff.pdfDRMI_withPRCff.png
Message ID: 14288     Entry time: Sat Nov 10 17:32:33 2018     In reply to: 14285     Reply to this: 14292
Author: gautam 
Type: Update 
Category: LSC 
Subject: Nulling MICH->PRCL and MICH->SRCL 

With the DRMI locked, I drove a line in MICH using the sensing matrix infrastructure. Then I looked at the error points of MICH, PRCL and SRCL. Initially, the sensing line oscillator output matrix for MICH was set to drive only the BS. Subsequently, I changed the --> PRM and --> SRM matrix elements until the line height in the PRCL and SRCL error signals was minimized (i.e. the change to PRCL and SRCL due to the BS moving, which is a geometric effect, is cancelled by applying the opposite actuation to the PRM/SRM respectively. Then I transferred these to the LSC output matrix (old numbers in brackets).

MICH--> PRM = -0.335 (-0.2655)

MICH--> SRM = -0.35 (+0.25)

I then measured the loop TFs - all 3 loops had UGFs around 100 Hz, coinciding with the peaks of the phase bubbles. I also ran some sensing lines and did a sensing matrix measurement, Attachment #1 - looks similar to what I have obtained in the past, although the relative angles between the DoFs makes no sense to me. I guess the AS55 demod phase can be tuned up a bit.

The demodulation was done offline - I mixed the time series of the actuator and sensor signals with a "local oscillator" cosine wave - but instead of using the entire 5 minute time series and low-passing the mixer output, I divvied up the data into 5 second chunks, windowed with a Tukey window, and have plotted the mean value of the resulting mixer output.

Unrelated to this work: I re-aligned the PMC on the PSL table, mostly in Pitch.

Attachment 1: sensMat_2018-11-10.pdf  53 kB  | Hide | Hide all
sensMat_2018-11-10.pdf
ELOG V3.1.3-